A grazes 10 sheeps for 3 weeks and B grazes 15 sheeps for 4 weeks in a field. How would they divide a rent of ` 60
Answer by richwmiller(17219)

10*3= 30 sheep weeks
15*4=60 sheep weeks
30/90 should pay 20/60=1/3
60/90 should pay 40/60=2/3